# SNF-ROM: Projection-based nonlinear reduced order modeling with smooth neural fields

### v1

@@ 6 Numerical experiments
$u_{0}(x,y;\mu)=\begin{cases}\mu\sin\left(2\pi x\right)\sin\left(2\pi y\right)&x,y\in[0,0.5]\\ 0&\text{otherwise}\end{cases}.$ (59)
The $x$ and $y$ components of $\boldsymbol{u}$ will be equal for all times as the problem is symmetric. Although the ROM models solve for both components, we restrict our analysis in Figure 11 and Figure 12 to the first component of the solution vector without loss of generality.
The solution at multiple time instances for different ROMs is shown in Figure 11. Similar to the 1D burgers test case, we observe that POD-ROM exhibits large oscillations and deviation from the FOM results at later times. All nonlinear model reduction approaches capture the shock well and exhibit much better accuracy than POD-ROM. A qualitative assessment of $\epsilon(\boldsymbol{x},t)$ indicates that CAE-ROM exhibits large errors near the shock front. SNFL-ROM and SNFW-ROM produce much smaller errors around the shock, as well as further away, with SNFW-ROM producing $\epsilon\leq 0.1\%$ errors everywhere at $t=0.5$.
This observation is more evident from results in Figure 12, which show ROM predictions along the diagonal slice $y=x$. We observe that POD-ROM exhibits a much higher error throughout the domain than nonlinear ROMs, which produce much more accurate results. The temporal evolution of error for different ROMs is shown in Figure 13. These results indicate that POD-ROM and CAE-ROM exhibit large errors that grow with time. Conversely, SNFL-ROM and SNFW-ROM experience a decrease in error at an early time, after which the error in SNFL-ROM rises steadily till the final time step, whereas the error in SNFW-ROM stabilizes to $\sim 0.1\%$. These results highlight the suitability of smooth neural fields, especially with weight regularization, to capture the physics of this 2D flow with a shock wave accurately and efficiently.
@@
Figure 15: 1D Kuromoto-Sivashinsky test case: variation of error in ROM predictions against time. Despite large oscillations in error values, the error in nonlinear ROMs remains below $1\%$ at all times. SNFW-ROM produces low error values at all times. Note that the $y$-axis in this figure has $\log$ scale, which may distort the oscillations in the error curves for the nonlinear ROMs.
A comparison of solutions using different ROMs at two time instances is shown in Figure 14. The solution grows from a Gaussian initial condition to a complex signal with varying magnitudes. As the problem is highly diffusive, all ROMs appear to give similar results that are in good agreement with FOM. To compare these approaches more closely, we analyze the temporal evolution of error defined in Eq. 51. The temporal evolution of model error for different ROMs is shown in Figure 15. POD-ROMs exhibit the highest error for all instances of time that are tested. CAE-ROMs exhibit a lower error value that remains consistent over time. SNFL-ROMs exhibit a much smaller error in the initial time window, which grows considerably with time and reaches the same magnitude as CAE-ROM. SNFW-ROM, on the other hand, exhibits low error at all times of approximately $0.2\%$. This result follows the observation from other test cases that weighted decay regularization yields consistently more accurate results. We observe large oscillations in the error curves for SNFL-ROM and SNFW-ROM in Figure 15 whose origin is not fully understood. However, despite these oscillations, the errors are consistently below those for POD-ROM and CAE-ROM indicating superior performance of proposed SNF-ROMs.
@@ 7 Conclusions and discussion
SNF-ROM is a machine learning-based nonlinear model order reduction method that predicts dynamics by combining neural field spatial representations from point-cloud datasets with Galerkin projections. Compared to existing nonlinear ROMs, SNF-ROM presents several advantages in the online stage of dynamics prediction. SNF-ROM constrains the learned ROM manifold to be smooth and regular, thus allowing for robust online stage evaluations with large time steps. By supporting AD-based spatial differencing, SNF-ROM drastically simplifies the computation of online dynamics. Numerical experiments reveal that the dynamics evaluation of SNF-ROM is accurate and that the proposed smoothing methodology is robust, thus eliminating the need for hyperparameter tuning. While offering the key advantages mentioned above, SNF-ROM consistently outperforms other tested ROM approaches across various PDE problems.
SNF-ROM is not without limitations. For one, the offline training cost is significantly greater than that for a linear POD-ROM. This higher cost is a significant hurdle in working with nonlinear ROMs, as large training times bottleneck prototyping and experimentation processes. Future work may consider using modern neural architectures that support fast training [39]. Another possible research direction could be to use transformer-based meta-learning approaches to learn network weights [40] quickly. Many nonlinear ROMs are unable to handle complex and changing boundary conditions; future work could be directed towards alleviating this limitation. Future work may also consider lowering the cost of the online stage by integrating novel hyper-reduction schemes within SNF-ROM. Further performance gains may be achieved by writing customized GPU kernels for the online stage and employing highly efficient Taylor mode AD [41, 42] instead of forward mode AD. Nonlinear reduced order modeling is an emerging technology in its nascent stage. This work has identified and eliminated critical shortcomings in existing approaches. Substantial enhancements must be made for its maturation and practical application.
@@ Acknowledgements
Appendix A CAE-ROM: Manifold projection
@@ Appendix B SNF-ROM: Manifold projection
However, our experiments in Section 6 indicate that $\Xi_{\varrho}(t,\boldsymbol{\mu})$ is typically not the local optimum of the system
@@ Appendix C SNF-ROM: Architecture and training details
We normalize the data set so that the spatial coordinates, the field values, and the times have zero mean and unit variance. Our training pipeline disaggregates simulation snapshots into tuples $(\boldsymbol{x},t,\boldsymbol{\mu},\boldsymbol{u}(\boldsymbol{x},\cdot))$ and stochastically composes the dataset $\mathcal{D}$ into $100$ equally-sized mini-batches. For SNFL-ROM, we use the Adam optimizer [45] for learning $\theta$, and $\varrho$. For SNFW-ROM, we use the AdamW optimizer [46] is employed as described in Section 4.2.2. While we formulate W regularization as an additional loss term, it can be easily implemented by modifying the weight-decay approach in an AdamW optimizer. Rather than computing the sensitivity of $L_{\text{Weight}}$ in Eq. 39 with respect to $\theta$ via backpropagation, the expressions can be written in closed form as
@@
with a base learning rate of $\eta=10^{-4}$. Additionally, we run a learning rate warmup cycle for $10$ epochs at $\eta=0.01$. Early stopping is imposed if the loss value does not decrease for $20\%$ of consecutive epochs at any learning rate.

### v2

@@ 6 Numerical experiments
$u_{0}(x,y;\mu)=\begin{cases}\mu\sin\left(2\pi x\right)\sin\left(2\pi y\right)&x,y\in[0,0.5]\\ 0&\text{otherwise}\end{cases}$ (60)
which is parameterized by a scalar $\mu\in\mathbb{R}_{+}$. The solution to this problem involves propagation of a wavefront that eventually forms a shock. The $x$ and $y$ components of $\boldsymbol{u}$ will be equal for all times as the problem is symmetric. Although the ROM models solve for both components, we restrict our analysis in Figure 12 and Figure 13 to the first component of the solution vector without loss of generality. We train the ROM models on FOM simulations with $\mu_{\text{train}}=\{0.900,0.933,0966,1.033,1.066,1.100\}$, and evaluate for $\mu_{\text{test}}=\{1.000\}$. This test enables the comparison of ROM approaches for parameter space interpolation. The results discussed hereafter are for the interpolation case $\mu=1.00$.
The solution at multiple time instances for different ROMs is shown in Figure 12. Similarly to the 1D burgers test case, we observe that POD-ROM exhibits large oscillations and deviation from the FOM results at later times. All nonlinear model reduction approaches capture the shock well and exhibit much better accuracy than POD-ROM. A qualitative assessment of $\epsilon(\boldsymbol{x},t)$ indicates that CAE-ROM exhibits large errors near the shock front. SNFL-ROM and SNFW-ROM produce much smaller errors around the shock, as well as further away, with SNFW-ROM producing $\epsilon\leq 0.1\%$ errors everywhere at $t=0.5$.
This observation is more evident from results in Figure 13, which show ROM predictions along the diagonal slice $y=x$. We observe that POD-ROM exhibits a much higher error throughout the domain than nonlinear ROMs, which produce much more accurate results. We also observe that CAE-ROM does not adequately capture the shock at $t=0.5$, whereas SNFL-ROM and SNFW-ROM are indistinguishable from the FOM solution. The temporal evolution of error for different ROMs is shown in Figure 14. These results indicate that POD-ROM and CAE-ROM exhibit large errors that grow with time. Conversely, SNFL-ROM and SNFW-ROM experience a decrease in error at an early time, after which the error in SNFW-ROM rises steadily till the final time step, whereas the error in SNFL-ROM stabilizes to $\sim 0.1\%$. These results highlight the suitability of smooth neural fields in capturing the physics of this 2D flow with a shock wave accurately and efficiently.
[…]
                         Number of hyper-reduction points (|$(\absolutevalue{X_{\text{proj}}})$
$16,384$ $4,096$ $1,024$ $256$ $64$
[…]
The first trend we note in Table 4 is that memory allocation decreases proportionally as we reduce |$\absolutevalue{X_{\text{proj}}}$ | and increase $\Delta t$. This is expected as reducing |$\absolutevalue{X_{\text{proj}}}$ | leads to evaluating $g_{\theta}$ on smaller batches, and increasing $\Delta t$ leads to fewer calculations altogether. We note that speed-up increases proportionally with $\Delta t$ for all |$\absolutevalue{X_{\text{proj}}}$ |. We also note a similar steady rise speed-up with a reduction in |$\absolutevalue{X_{\text{proj}}}$ | till |$\absolutevalue{X_{\text{proj}}}=256$. Thereafter, we observe no substantial increase in speed-up as the wall-time is dominated by latency of CPU to GPU communication. The maximum speed-up achieved in this experiment is 199×199$\times$199 × against the FOM calculation.
In Table 4, we note that the relative error increases as we decrease |$\absolutevalue{X_{\text{proj}}}$ |. This is expected as the dynamics associated with the omitted collocation points is not available to guide the time-integration of the reduced system. We also note that, as expected, the relative error increases steadily with $\Delta t$. Both trends in relative error are thwarted by the case with $\Delta t=10\Delta t_{0}$ and $X_{\text{proj}}=64$, which has a relative error of $0.369\%$. This might be due to fortuitous cancellation of error introduced by hyper-reduction and time-integration, although further investigation is needed to ascertain the reason for this phenomenon.
@@
Figure 16: 1D Kuromoto-Sivashinsky test case: variation of error in ROM predictions against time. Despite large oscillations in error values, the error in nonlinear ROMs remains below $1\%$ at all times. SNFL-ROM and SNFW-ROM produce low error values at all times. Note that the $y$-axis in this figure has $\log$ scale, which may distort the oscillations in the error curves for the nonlinear ROMs.
A comparison of solutions using different ROMs at two time instances is shown in Figure 15. The solution grows from a Gaussian initial condition to a complex signal with varying magnitudes. As the problem is highly diffusive, all ROMs appear to give similar results that are in good agreement with FOM. To compare these approaches more closely, we analyze the temporal evolution of the error defined in Eq. 52. The temporal evolution of model error for different ROMs is shown in Figure 16 which presents the error evolution of ROMs from a dynamics evaluation with $\Delta t=\Delta t_{0}$ and a dynamics evaluation with $\Delta t=10\Delta t_{0}$. POD-ROMs exhibit the highest error for all instances of time that are tested with both $\Delta t$. CAE-ROMs exhibit a lower error value that remains consistent over time for $\Delta t=\Delta t_{0}$. For $\Delta t=10\Delta t$, however, the error in POD-ROM rises dramatically with time. SNF-ROMs exhibit a much smaller error in the initial time window, which does not grow considerably over time. We observe large oscillations in the error curves for SNFL-ROM and SNFW-ROM in Figure 16 whose origin is not fully understood. However, despite these oscillations, the errors are consistently below those for POD-ROM and CAE-ROM indicating superior performance of proposed SNF-ROMs. Low errors in this numerical experiment indicate that SNF-ROM can adequately capture the dynamics of PDE problems with higher-order derivatives.
@@ 7 Conclusions and discussion
SNF-ROM is a machine learning-based nonlinear model order reduction method that predicts dynamics by combining neural field spatial representations from point-cloud datasets with Galerkin projections. Compared to existing nonlinear ROMs, SNF-ROM presents several advantages in the online stage of dynamics prediction. SNF-ROM constrains the learned ROM manifold to be smooth and regular, thus allowing for robust online stage evaluations with large time steps. By supporting AD-based spatial differencing, SNF-ROM drastically simplifies the computation of online dynamics. Numerical experiments reveal that the dynamics evaluation of SNF-ROM is accurate and that the proposed smoothing methodology is robust, thus eliminating the need for significant hyperparameter tuning. With a combination of hyper-reduction and larger time-steps, SNF-ROM faster than the full-order model by up to 199×199$\times$199 ×. While offering the key advantages mentioned above, SNF-ROM consistently outperforms other tested ROM approaches across various PDE problems.
SNF-ROM is not without limitations. For one, the offline training cost is significantly greater than that for a linear POD-ROM. This higher cost is a significant hurdle in working with nonlinear ROMs, as large training times bottleneck prototyping and experimentation processes. Future work may consider using modern neural architectures that support fast training [39]. Another possible research direction could be to use transformer-based meta-learning approaches to learn network weights [40] quickly. Many nonlinear ROMs are unable to handle complex and changing boundary conditions; future work could be directed towards alleviating this limitation. Further performance gains may be achieved by writing customized GPU kernels for the online stage and employing highly efficient Taylor mode AD [41, 42] instead of forward mode AD. Nonlinear reduced order modeling is an emerging technology in its nascent stage. This work has identified and eliminated critical shortcomings in existing approaches. Substantial enhancements must be made for its maturation and practical application.
@@ Acknowledgements
Appendix A Manifold projection
A.1 CAE-ROM
@@ A.2 SNF-ROM
However, our experiments in Section 6 indicate that $\Xi_{\varrho}(t,\boldsymbol{\mu})$ may not be the local optimum of the system
@@
Appendix B Neural field regularization
Here, we discuss the effectiveness of Lipschitz regularization and weight regularization at smoothing neural field representations. The purpose of this step is to ensure that spatial derivatives of neural field representations adequately capture the derivatives of the true signal. A typical neural field representation is not guaranteed to smoothly interpolate the training data. As such, partial derivatives of neural field ROMs are riddled with numerical artifacts, making them unusable in downstream applications such as time-evolution. In this appendix, we assess the effectiveness of the proposed regularization strategies in Section 4.2 at curbing numerical artifacts in spatial derivatives of a neural field representation.
In Figure 17, we train a set of MLPs to regress a simple $1D$ function with different regularization strategies. We fix the width of the hidden layer of each MLP to be $64$ and adopt the training strategy and architecture details of $g_{\theta}$ described in C. Spatial derivatives of the neural network representations are then computed exactly with automatic differentiation and compared with the true derivatives of the signal $u(x)$ in Figure 17.
Table 5 quantifies the mismatch between the derivatives of the neural field representations and the derivatives of the true signal by computing the relative mean-squared error. We note that all neural fields accurately capture the ground truth function. However, when no regularization is applied, the first and second derivatives differ from those of the true signal. This may lead to numerical instabilities in the dynamical evaluation in Section 5. In contrast, both Lipschitz regularization and weight regularization adequately capture the derivative of the signal. This is evident in the successful numerical experiment in Section 6.5 where we perform dynamics evaluation on the fourth-order KS problem.
Figure 17: We compare the effectiveness of Lipschitz regularization and weight regularization at smoothing neural field representations. We train a set of neural networks to regress the function $u(x)=(x-\dfrac{\pi}{2})\sin(x)\exp(-x^{2})$ with different regularization strategies (top). The first (middle) and second (bottom) derivatives of the neural network representations are computed exactly in automatic differentiation. When no regularization (orange line) is applied, the first and second derivatives of the neural field differ from those of the true signal. Comparatively, Lipschitz regularization and weight regularization significantly smooth out the noise in spatial derivatives of the neural field representation.
Table 5: We compare the effectiveness of Lipschitz regularization and weight regularization at smoothing neural field representations. We compare the derivatives of the signal $u(x)=(x-\dfrac{\pi}{2})\sin(x)\exp(-x^{2})$ and those of neural field representations trained to regress $u(x)$ with different regularization strategies. The mismatch is quantified with the relative mean squared error.
Regularization approach $u(x)$ $u^{\prime}(x)$ $u^{\prime\prime}(x)$
No regularization $2.75\times 10^{-6}$ $7.06\times 10^{-2}$ $70.1$
Lipschitz regularization ($\alpha=5\times 10^{-5}$) $3.68\times 10^{-6}$ $5.28\times 10^{-4}$ $9.72\times 10^{-2}$
Weight regularization ($\gamma=5\times 10^{-2}$) $2.10\times 10^{-6}$ $1.01\times 10^{-4}$ $1.76\times 10^{-2}$
@@ Appendix C SNF-ROM: Architecture and training details
We normalize the data set so that the spatial coordinates, the field values, and the times have zero mean and unit variance. Our training pipeline disaggregates simulation snapshots into tuples $(\boldsymbol{x},t,\boldsymbol{\mu},\boldsymbol{u}(\boldsymbol{x},\cdot))$ and stochastically composes the dataset $\mathcal{D}$ into $100$ equally-sized mini-batches. For SNFL-ROM, we use the Adam optimizer [45] for learning $\theta$, and $\varrho$. For SNFW-ROM, we use the AdamW optimizer [46] is employed as described in Section 4.2.2. While we formulate weight regularization as an additional loss term, it can be easily implemented by modifying the weight-decay approach in an AdamW optimizer. Rather than computing the sensitivity of $L_{\text{Weight}}$ in Eq. 39 with respect to $\theta$ via backpropagation, the expressions can be written in closed form as
@@
with a base learning rate of $\eta=10^{-4}$. For the 2D Burgers problem, we reduce the number of epochs from $200$ to $30$ per learning rate. This is done to reduce training time. Additionally, we run a learning rate warmup cycle for $10$ epochs at $\eta=0.01$. Early stopping is imposed if the loss value does not decrease for $20\%$ of consecutive epochs at any learning rate. When training is complete, we find that all models have $\overline{L}_{\text{data}}(\theta,\varrho;\mathcal{D}_{\text{train}})\sim 10^{-6}$ or lower.